OpenStudy (anonymous):

You throw a stone vertically upward with an initial speed of 6.0m/s from a third-story office window. If the window is 10m above the ground, find the time the stone is in flight. Find the speed of the stone just before it hits the ground. Idk why but I'm having trouble with this

OpenStudy (anonymous):

You're either forgetting the initial displacement, or you're getting the signs wrong so up and down are being mixed up. You use the displacement equation with a non-zero initial position, and you solve for when the height is zero. Remember that up is positive. Dfinal = Dinitial + Vinitial * t + (1/2)*a*t^2 0 = 10 + 6*t - (1/2)*9.8*t^2 Solve for t. If you use the quadratic formula, you'll get two possible answers. You want the positive one. FYI, it's somewhat more than two seconds, but I'll let you do the arithmetic. Once you have t from the first part, you can calculate the velocity when it hits the ground using Vfinal = Vinitial + a*t Vfinal = 6 - 9.8*t The speed is the absolute value of this, somewhat greater than 15 m/s downwards. But again, I'll let you do the arithmetic. :)